OpenStudy (anonymous):

What is the discontinuity of the function f(x) = the quantity of x squared minus 4 x minus 12, all over x plus 2? (−6, 0) (6, 0) (−2, −8) (2, −4)

OpenStudy (mathmale):

Please translate that statement into a mathematical equation. This is an important basic skill. \[f(x)=\frac{ x^2-4x-12 }{ x-2 }\]

OpenStudy (mathmale):

Now look carefully at the denominator. Which x value would make the denom. = 0?

OpenStudy (mathmale):

Question for you: How do you specify a "discontinuity?"

OpenStudy (anonymous):

x+2 would make the denom = 0

OpenStudy (mathmale):

So, you're saying, "If we add together x+2, that would make the denom. = 0" Is that true? Aren't you supposed to be finding the value of x that makes the denom. = 0?

OpenStudy (anonymous):

my teacher didnt explain how to do this type of problem very well

OpenStudy (anonymous):

so i have little knowledge on how to do this

OpenStudy (mathmale):

To answer the question, just let the whole denom. = 0. Solve for x. If x-2=0, adding 2 to both sides of the eq'n will give you a value for x. What is that value? It is there that your function has a discontinuity, since division by zero is not allowed.

OpenStudy (mathmale):

sorry you're not happy with your teacher. You could, however, ask questions here on OpenStudy, and thus possibly learn what you did not learn in class.